OpenStudy (anonymous):

What is the slope of the line passing through the points (0, 4) and (–8, –1)?

OpenStudy (nurali):

m=y2-y1/x2-x1

OpenStudy (nurali):

(x1,y1)=(0,4) (x2,y2)=(-8,-1)

OpenStudy (anonymous):

a) -8/5 b)-5/8 c)5/8 d) 8/5

OpenStudy (nurali):

m=-1-4/-8-0

OpenStudy (nurali):

m=?

OpenStudy (nurali):

reply

OpenStudy (anonymous):

im trying to understand what u mean srry

OpenStudy (anonymous):

srry didnt mean to send that

OpenStudy (nurali):

give me answer.

OpenStudy (anonymous):

c

OpenStudy (nurali):

Correct.

OpenStudy (anonymous):

really?

OpenStudy (nurali):

Yes

OpenStudy (anonymous):

ok thanks

OpenStudy (nurali):

My Pleasure.